Dess:

Suppose an object is thrown upward with an initial velocity of 52 feet per second from a height of 125 feet. The height of the object t seconds after it is thrown is given by h(t)=-16t^2+52t+125. Find the average velocity in the first three seconds after the object is thrown.

SkyVoltage43:

give me a minute im trying to work out all the math

Dess:

ok

SkyVoltage43:

do you have options

Dess:

no it'd an assignment where I have to work it out

SkyVoltage43:

ok

SkyVoltage43:

do you know how to work it out you just dont know the answer?

Dess:

I don't know either. I'll just ask my teacher in class

SkyVoltage43:

ok well the answer is 18

Dess:

how did you get 18?

SkyVoltage43:

idk i just looked at it and got 18, I did mental math some how and got 18

Dess:

I don't think that's right, because it has a formula.

SkyVoltage43:

the average velocity in the first three seconds after the object is thrown Would be 18

SkyVoltage43:

have your teacher help you but im sure my answer is right

celticcat:

The velocity at time t s given by the derivative of h(t) V = -32t + 52

celticcat:

so at time 0 v = 52 and at time 3 , V = -(32)3 + 52 = - 44 ( at this time the object is descending)

Dess:

Thanks for explaining, is there another step I need?

celticcat:

they are asking for the average velocity Hmmm Im not sure if we can use these 2 results to get that...

SkyVoltage43:

its 18

celticcat:

How did you get that?

SkyVoltage43:

just trust me

Dess:

I'm confused now

sillybilly123:

\(h(0) = 125\) \(h(3) = 137\) Net ***Displacement*** over the 3 second period \(= 137 - 125 = 12\) Ave Velocity over that period \(= \dfrac{12}{3} = 4\)

sillybilly123:

units....ft/ sec

Dess:

How did you get the 137?

Falconmaster:

starfish comand

sillybilly123:

it's h(3), the height at t = 3

celticcat:

yea thats the correct method of course. I guess you cant average velocities if an object is accelerating.

celticcat:

the displacement from the original starting point is 12ft and velocity = displacement / time

Dess:

1 attachment
SkyVoltage43:

this is starfish command what do you need bubblez?

SkyVoltage43:

@Falconmaster

Dess:

@sillybilly123

celticcat:

what do you want to know about that formula?

Dess:

Is formula that @sillybilly123 use?

Dess:

Because this formula is used for this assignement

celticcat:

basically yes f(b) and f(a) will be the values of h(t) at times 3 and 0 and b-a = 3 - 0

celticcat:

137 - 125


3 - 0

celticcat:

follow?

celticcat:

average velocity = [ (h3) - h(0) ] / (3-0) = (137 - 125) / 3 = 12/3 = 4

Dess:

so the answer is 4?

Dess:

@celticcat

celticcat:

yes

Falconmaster:

starfish command

Dess:

Yes for the answer being 4?

Dess:

Sorry I'm a little slow.

ldenault:

i have no freaking clue because im only in middle school

ldenault:

you can just look at my beautiful PFP